# MetaMath: Bootstrap Your Own Mathematical Questions for Large Language Models

### v1

@@ 3.4 Finetuning Objective Functions
We merge all the augmented data, including answer-augmented data and bootstrapped questions (Rephrasing, Self-Verification, FOBAR) as:
$\mathcal{D}_{\text{MetaMathQA}}=\mathcal{D}_{\text{AnsAug}}\cup\mathcal{D}_{\text{rephrase}}\cup\mathcal{D}_{\text{SV}}\cup\mathcal{D}_{\text{FOBAR}}.$ (5)
We finetune a LLM model (parameterized by ${\bm{\theta}}$) on $\mathcal{D}_{\text{MetaMathQA}}$ to obtain the MetaMath model by maximizing the log likelihood of the reasoning path conditioned on the question, i.e.,
$\mathcal{L}({\bm{\theta}})=\sum_{(q,r,a)\in\mathcal{D}_{\text{MetaMathQA}}}\log\mathbb{P}(r\mid q;{\bm{\theta}}).$ (6)
Although we only consider LLaMA-2 here, MetaMathQA can also be used to finetune other LLMs.
@@ 4.1 Experimental Setup
Dataset AnsAug Rephrased SV FOBAR Overall
@@
Datasets. We use two popular mathematical reasoning benchmarks: (i) GSM8K [10] is a dataset consists of high quality grade school math problems, containing 7,473 training samples and 1,319 testing samples; and (ii) MATH [19] dataset consists of high school math competition problems that span seven subjects including Prealgebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra, and Precalculus. It contains 7,500 and 5,000 samples for training and testing, respectively. Questions in GSM8K [10] take between 2 and 8 steps to reach the answer, while MATH is much more challenging.
Models. We use the current state-of-the-art open-source model LLaMA-2 [57], including three different parameter sizes: 7B, 13B, and 70B, as base model for fine-tuning. GPT-3.5-Turbo is used for rephrasing questions as well as generating answers in all four augmentations. The LLaMA-2-7B and LLaMA-2-13B are trained by fully fine-tuning. LLaMA-2-70B is fine-tuned by QLoRA [12] for computational efficiency. More experimental details can be seen in Appendix A.2.
Baselines. The proposed methods are compared with (i) closed-source models such as GPT-3.5-Turbo [43], PaLM [9]; (ii) open-source models such as LLaMA-1 [56], LLaMA-2 [57]; (iii) Supervised Fine-Tuning (SFT), which uses the training set of the original GSM8K or MATH datasets for fine-tuning; (iv) Rejection sampling Fine-Tuning (RFT) [64] generates and collects correct reasoning paths as augmented data for fine-tuning; (v) WizardMath [34] which generates samples and trains two reward models using ChatGPT  [^2] https://openai.com/ to select samples for fine-tuning.
Diversity Gain. We use the diversity gain [4] to measure to what extent a new dataset added to a basic dataset can improve the overall data diversity. For a base dataset $\mathcal{D}_{base}=\{x_{i}=(q_{i},r_{i},a_{i})\}_{i=1}^{N}$ with $N$ samples, and a new dataset $\mathcal{D}_{new}=\{x_{i}=(q_{i},r_{i},a_{i})\}_{i=1}^{M}$ with M samples, the diversity gain is defined as: $\mathcal{D}_{new}$ relative to $\mathcal{D}_{base}$ as: $d_{gain}=\frac{1}{M}\sum_{x_{i}\in\mathcal{D}_{new}}\min_{x_{j}\in\mathcal{D}_{base}}(\|f(x_{i})-f(x_{j})\|_{2}^{2})$, where $f$ is the feature extractor and we use the OpenAI Embedding API text-embedding-ada-002 for feature extraction. For Figure 2, we change the data size of base data and select a fixed set of 20K new data points that the model has not encountered to form $\mathcal{D}_{new}$.
@@
${}^{\ddagger}$ 70B 82.3 26.0
Table 2: Comparison of testing accuracy to existing LLMs on GSM8K and MATH. ${}^{\ddagger}$ue to the computing resource limitation, we finetune MetaMath-70B using QLoRA [12].
@@ 4.2 Results on GSM8K and MATH
Table 1 illustrates the detailed description of our MetaMathQA collection and Table 2 shows the testing accuracy on GSM8K and MATH. As can be seen, for open-source models with 1-10B parameters, MetaMath achieves the state-of-the-art performance. Compared to the previous best LLM, MetaMath achieves a large improvement of 11.5% on GSM8K and 8.7% on MATH in testing accuracy, showing that finetuning on our MetaMathQA data is effective.
@@ 4.3 Effect of Augmentations
In this section, we conduct experiments to study the effect of augmentations in MetaMath. We first fine-tune the LLaMA-2-7B model on augmented GSM8K data, and test the fine-tuned model on GSM8K and MATH. Table 3 shows the testing accuracy of different combinations of augmentations. As can be seen, on GSM8K, the MetaMath models trained on answer augmentation (AnsAug) or rephrasing augmentation achieve much higher accuracy than SFT, which is only trained on the training set. Combing answer augmentation and rephrasing augmentation data for fine-tuning leads to a slightly higher accuracy, which is further improved by about 4% through merging the FOBAR and SV augmentation data. As for MATH, MetaMath trained on augmented GSM8K data performs better than SFT, suggesting its effectiveness in generalizing to unseen mathematical tasks. We also conduct an experiment by fine-tuning LLaMA-2-7B on the augmented MATH data then evaluate the model on GSM8K and MATH. Table 3 shows the testing accuracy. Again, MetaMath trained on AnsAug or rephrasing augmentation data performs much better than SFT. Furthermore, merging all augmented data together for fine-tuning is better than merging AnsAug and rephrasing augmentation data, demonstrating the effectiveness of SV and FOBAR augmentation data in improving mathematical reasoning ability. Moreover, for the unseen GSM8K task, MetaMath trained on augmented MATH data is significantly better than SFT (+15%).
@@
4.4 Reasoning Paths with Incorrect Answer Can Also Be Useful
Method #samples accuracy
SFT [57] 7, 473 41.6
Incorrect Answers 7, 473 43.6
Correct Answers 7, 473 47.3
Table 4: Testing accuracy on GSM8K of LLaMA-2-7B trained on answer augmented data with correct/incorrect answer.
We conduct experiments on GSM8K using LLaMA-2-7B to study whether the answer augmentation samples with incorrect answers are helpful for finetuning the LLM. We randomly choose 7,473 reasoning paths with incorrect answers from the generated answers, and we ensure that the size is the same as that of the original training set. From Table 4, we observe that the model finetuned on the augmented data with incorrect answers is actually better than SFT, which is counter-intuitive. We hypothesize that although the final answer is incorrect, some intermediate reasoning steps are correct (see Example LABEL:exmp:incorrect). These reasoning steps can still be useful supervision signals. Our results are also aligned with Lightman et al. [31], where they also discover the importance of intermediate process supervision for reasoning.
{exmp}
A Reasoning Path with Incorrect AnswerincorrectQuestion: Daisy and Rose were enjoying their backyard pool with their dogs. If there are 24 legs/paws in the pool, how many dogs do Daisy and Rose have? (Ground-truth answer is $5$)
Generated Answer: Each dog has 4 legs, so if there are 24 legs in total, there are 24/4 = 6 dogs. Therefore, Daisy and Rose have 6 dogs. The answer is: 6.
Figure 3: Lower perplexity of MetaMathQA.
@@ 4.5 Discussion from a Perplexity Perspective
According to the Superficial Alignment Hypothesis proposed by  Zhou et al. [67], the capability of a model is rooted in pretraining, and data from downstream tasks acts to activate the inherent ability of LLMs that has been learned during pretraining. There are two important questions that arise from such a hypothesis: (i) what kind of data is most effective at activating possible latent knowledge, and (ii) why is one dataset better than another at such activation? Our empirical results suggest that, in the mathematical tasks we consider, our MetaMathQA dataset may serve as a superior activator of the mathematical knowledge. Yet, why MetaMath yields superior performance than training on the data of correct answer-only or GSM8K CoT is unclear. We speculate that perhaps it is the simplicity of the data that matters. As shown in Figure 4, we compute the perplexity [59, 37] for the under-finetuned LLaMA-2-7B model, in terms of answer-only data, GSM8K CoT, and the subsections of MetaMathQA data. The perplexity of MetaMathQA is significantly lower than the other two datasets. This highlights its inherently easy-to-learn nature, which may be more conducive to eliciting bolstered problem-solving abilities from an LLM. This is also aligned with the findings with TinyStories [14], where short and easy story data can help LLMs generate content fluently.
@@ 4.6 Discussion from a Diversity perspective
4.7 More Data Can Actually Hurt the Performance
Figure 6: Testing accuracy on the questions with or without the “IF” statement.
Figure 6: Testing accuracy on the questions with or without the “IF” statement.
@@ 4.7 More Data Can Actually Hurt the Performance
{exmp}
Case study in test set of GSM8Kcase-studyQuestion: Darrell and Allen’s ages are in the ratio of 7:11, If their total age now is 162, calculate Allen’s age 10 years from now. (Ground-truth answer is $109$)
SFT Answer: The total ratio representing their ages is 7+11=<<much-less-than$<$ <7+11=18>>much-greater-than$>$ >18. The fraction representing Darrell’s age is 7/18, and since their total age now is 162, Darrell’s age is 7/18*162 = <<much-less-than$<$ <7/18*162=60>>much-greater-than$>$ >60. The fraction representing Allen’s age is 11/18, and since their total age now is 162, Allen’s age is 11/18*162 = <<much-less-than$<$ <11/18*162=84>>much-greater-than$>$ >84. Ten years from now, Allen’s age will be 84 + 10 = <<much-less-than$<$ <84+10=94>>much-greater-than$>$ >94.
RFT Answer: The total ratio representing Darrell and Allen’s ages is 7+11 = <<much-less-than$<$ <7+11=18>>much-greater-than$>$ >18. The fraction representing Darrell’s age from the total ratio is 7/18, and since their total age is 162, Darrell is 7/18*162 = 100 years old. Allen’s age, from the total ratio, is 11/18, and since Darrell is 100, Allen is 11/18*109 = <<much-less-than$<$ <11/18*100=66>>much-greater-than$>$ >66 years old. Ten years from now, Allen will be 66 + 10 = 76 years old.
Our Answer: The ratio of Darrell’s age to Allen’s age is 7:11. Let’s assume Darrell’s age is 7x and Allen’s age is 11x. The total age of Darrell and Allen is 7x + 11x = 18x. We are given that the total age is 162, so 18x = 162. Dividing both sides by 18, we get x = 9. Therefore, Allen’s age is 11x = 11 * 9 = 99. 10 years from now, Allen’s age will be 99 + 10 = 109.
To understand the benefit brought by the proposed bootstrapping method, we conducted a case study on GSM8K. We evaluate different 7B mathematical LLMs on two groups in the test set: questions with “IF” and questions without “IF”. As can be seen in Figure 7, existing methods (SFT, RFT, WizardMath) have a significant accuracy gap between these two groups. In comparison, MetaMath has a balanced performance, and the performance on two separate sets is better than other models. One specific question is illustrated in Example LABEL:exmp:case-study. Given a question containing “IF”, different from SFT and RFT, our MetaMath generates the correct answer due to proposed backward bootstrapping.
@@ 5 Concluding Remarks
In this paper, we focus on improving the mathematical problem-solving abilities of open-source LLMs. By bootstrapping mathematical questions on GSM8K and MATH, we present a high-quality and diverse dataset MetaMathQA, involving forward reasoning and backward reasoning samples. Our family of LLMs fine-tuned on MetaMathQA, called MetaMath, have achieved state-of-the-art on mathematical benchmarks among all open-source LLMs. Remarkably, MetaMath-7B reaches $66.4\%$ on GSM8K and $19.4\%$ on MATH, surpassing previous open-source LLMs by a significant margin. Our work further emphasizes the importance of the characteristics of the training data on boosting LLM problem-solving capabilities.
Our work, despite being simple yet effective, is by no means perfect. We explore a few possible question bootstrapping strategies and gain a significant performance gain. What kind of augmented data can best improve the mathematical problem-solving ability of LLM and why the augmented data actually help are left as open problems. We expect that more future endeavors can be spent in exploring the intrinsic mechanism behind question augmentation for LLMs.
@@ A.1 Rephrasing Prompts
{exmp}
Prompt for Rephrasing GSM8K QuestionsRephrase-prompt-allYou are an AI assistant to help me rephrase questions. Follow the given examples.
Question: Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left?
Rephrase the above question: What is the amount of money that Olivia has left after purchasing five bagels for \$3 each, if she initially had \$23?
Question: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday?
Rephrase the above question: After losing 23 golf balls on Tuesday and an additional 2 on Wednesday, how many golf balls does Michael have left if he initially had 58 golf balls?
Question: Angelo and Melanie want to plan how many hours over the next week they should study together for their test next week. They have 2 chapters of their textbook to study and 4 worksheets to memorize. They figure out that they should dedicate 3 hours to each chapter of their textbook and 1.5 hours for each worksheet. If they plan to study no more than 4 hours each day, how many days should they plan to study total over the next week if they take a 10-minute break every hour, include 3 10-minute snack breaks each day, and 30 minutes for lunch each day? Rephrase the above question: Angelo and Melanie need to study 2 chapters in their textbook and 4 worksheets for their upcoming test. They have planned to dedicate 3 hours for each chapter and 1.5 hours for each worksheet. They can study for a maximum of 4 hours each day, taking into account 10-minute breaks every hour, 3 10-minute snack breaks per day, and 30 minutes for lunch. How many days do they need to study in total over the next week to complete their study plan?
Question: Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total?
Rephrase the above question: If Leah had 32 chocolates and her sister had 42, and they both consumed 35 chocolates, what is the total number of chocolates that they have left?
Question: There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room?
Rephrase the above question: If there were initially nine computers in the server room and five more computers were added each day from Monday to Thursday, what is the current total number of computers in the server room?
Question: Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny?
Rephrase the above question: If Jason initially had 20 lollipops and now has 12 after giving some to Denny, how many lollipops did he give to Denny?
Question: Sam bought a dozen boxes, each with 30 highlighter pens inside, for \$10 each box. He rearranged five of these boxes into packages of six highlighters each and sold them for \$3 per package. He sold the rest of the highlighters separately at the rate of three pens for \$2. How much profit did he make in total, in dollars?
Rephrase the above question: Sam purchased 12 boxes, each containing 30 highlighter pens, at \$10 per box. He repackaged five of these boxes into sets of six highlighters and sold them for \$3 per set. He sold the remaining highlighters individually at a rate of three pens for \$2. What is the total profit he made in dollars?
Question: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today?
Rephrase the above question: If there were initially 15 trees in the grove and the grove workers are planning to plant more trees today, resulting in a total of 21 trees, how many trees did the workers plant today?
Question: {Q}
Rephrase the above question:
A.2 Experimental Details

### v4

@@ 3.4 Finetuning Objective Functions
We merge all the augmented data, including answer-augmented data and bootstrapped questions (Rephrasing, Self-Verification, FOBAR) as ${\mathcal{D}_{\text{MetaMathQA}}=\mathcal{D}_{\text{AnsAug}}\cup\mathcal{D}_{\text{rephrase}}\cup\mathcal{D}_{\text{SV}}\cup\mathcal{D}_{\text{FOBAR}}}$. We finetune a LLM model (parameterized by ${\bm{\theta}}$) on $\mathcal{D}_{\text{MetaMathQA}}$ to obtain the MetaMath model by maximizing the log likelihood of the reasoning path conditioned on the question, i.e., ${\mathcal{L}({\bm{\theta}})=\sum_{(q,r,a)\in\mathcal{D}_{\text{MetaMathQA}}}\log\mathbb{P}(r\mid q;{\bm{\theta}}).}$ Although we only consider LLaMA-2 here, MetaMathQA can also be used to finetune other LLMs.
@@ 4.1 Experimental Setup
Dataset AnsAug Rephrasing SV FOBAR Overall
@@
Datasets. We use two popular mathematical reasoning benchmarks: (i) GSM8K [13] is a dataset consisting of high-quality grade school math problems, containing 7,473 training samples and 1,319 testing samples; and (ii) MATH [23] dataset consists of high school math competition problems that span seven subjects including Prealgebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra, and Precalculus. It contains 7,500 and 5,000 samples for training and testing, respectively. Questions in GSM8K [13] take between 2 and 8 steps to reach the answer, while MATH is much more challenging.
@@
MetaMath 70B 82.3 26.6
Table 2: Comparison of testing accuracy to existing LLMs on GSM8K and MATH. Due to the computing resource limitation, we finetune MetaMath-70B using QLoRA [15].
Models. We use the current state-of-the-art open-source model LLaMA-2 [70], including three different parameter sizes: 7B, 13B, and 70B, as the base model for fine-tuning. GPT-3.5-Turbo is used for rephrasing questions as well as generating answers in all four augmentations, where the temperature is set to 0.7 as in [74]. The LLaMA-2-7B and LLaMA-2-13B are trained by fully fine-tuning. LLaMA-2-70B is finetuned by QLoRA [15] for computational efficiency. More experimental details can be seen in Appendix B.
Baselines. The proposed methods are compared with (i) closed-source models such as GPT-3.5-Turbo [53], PaLM [12]; (ii) open-source models such as LLaMA-1 [69], LLaMA-2 [70]; (iii) Supervised Fine-Tuning (SFT), which uses the training set of the original GSM8K or MATH datasets; (iv) Rejection sampling Fine-Tuning (RFT) [79] generates and collects correct reasoning paths as augmented data for fine-tuning; (v) WizardMath [43] which generates samples and trains two reward models using ChatGPT  [^1] https://openai.com/ to select samples for fine-tuning.
Diversity Gain. We use the diversity gain [6] to measure to what extent a new dataset added to a basic dataset can improve the overall data diversity. For a base dataset $\mathcal{D}_{\text{base}}=\{x_{i}=(q_{i},r_{i},a_{i})\}_{i=1}^{N}$ with $N$ samples, and a new dataset $\mathcal{D}_{\text{new}}=\{x_{i}=(q_{i},r_{i},a_{i})\}_{i=1}^{M}$ with M samples, the diversity gain is defined as: $\mathcal{D}_{\text{new}}$ relative to $\mathcal{D}_{\text{base}}$ as: $d_{\text{gain}}=\frac{1}{M}\sum_{x_{i}\in\mathcal{D}_{\text{new}}}\min_{x_{j}\in\mathcal{D}_{\text{base}}}(\|f(x_{i})-f(x_{j})\|_{2}^{2})$, where $f$ is the feature extractor and we use the OpenAI Embedding API text-embedding-ada-002 for feature extraction. For Figure 2, we change the data size of base data and select a fixed set of 20K new data points that the model has not encountered to form $\mathcal{D}_{\text{new}}$.
@@ 4.2 Results on GSM8K and MATH
Table 1 illustrates the detailed description of our MetaMathQA collection and Table 2 shows the testing accuracy on GSM8K and MATH. As can be seen, for open-source models with 1-10B parameters, MetaMath achieves the state-of-the-art performance. Compared to the previous best LLM, MetaMath achieves a large improvement of 11.6% on GSM8K and 9.1% on MATH in testing accuracy, showing that finetuning on our MetaMathQA data is effective.
@@ 4.3 Effect of Augmentations
In this section, we conduct experiments to study the effect of augmentations in MetaMath. We first finetune the LLaMA-2-7B model on augmented GSM8K (MetaMath-GSM8K) data, and test the finetuned model on GSM8K and MATH. Table 3 shows the testing accuracy of different combinations of augmentations, where we mix all augmented data together for each model. As can be seen, on GSM8K, the models trained on answer augmentation (AnsAug) or rephrasing augmentation achieve much higher accuracy than SFT, which is only trained on the training set. Combing answer augmentation and rephrasing augmentation data for fine-tuning leads to a slightly higher accuracy, which is further improved by about 4% through merging the FOBAR and SV augmentation data. As for MATH, MetaMath trained only on MetaMahQA-GSM8K data performs better than SFT, suggesting its effectiveness in generalizing to unseen mathematical tasks.
We also conduct an experiment by fine-tuning LLaMA-2-7B on the augmented MATH (MetaMathQA-MATH) data then evaluate the model on GSM8K and MATH. Table 3 shows the testing accuracy. Again, MetaMath trained on AnsAug or rephrasing augmentation data performs much better than SFT. Furthermore, merging all augmented data together for fine-tuning is better than merging AnsAug and rephrasing augmentation data, demonstrating the effectiveness of SV and FOBAR augmentation data in improving mathematical reasoning ability. Moreover, for the unseen GSM8K task, MetaMath trained on MetaMathQA-MATH data is significantly better than SFT (+20%).
@@ 4.4 Discussion from a Perplexity Perspective
According to the Superficial Alignment Hypothesis proposed by  Zhou et al. [83], the capability of a model is rooted in pretraining, and data from downstream tasks acts to activate the inherent ability of LLMs that has been learned during pretraining. There are two important questions that arise from such a hypothesis: (i) what kind of data is most effective at activating possible latent knowledge, and (ii) why is one dataset better than another at such activation? Our empirical results suggest that, in the mathematical tasks we consider, our MetaMathQA dataset may serve as a superior activator of mathematical knowledge. Yet, why MetaMath yields superior performance than training on the data of correct answer-only or GSM8K CoT is unclear. We speculate that perhaps it is the simplicity of the data that matters. As shown in Figure 4, we compute the perplexity [72, 46] for the under-finetuned LLaMA-2-7B model, in terms of answer-only data, GSM8K CoT, and the subsections of MetaMathQA data. The perplexity of MetaMathQA is significantly lower than the other two datasets. This highlights its inherently easy-to-learn nature, which may be more conducive to eliciting bolstered problem-solving abilities from an LLM. This is also aligned with the findings with TinyStories [18], where short and easy story data can help LLMs generate content fluently.
@@ 4.5 Discussion from a Diversity perspective
4.6 Evaluating the Reversal Mathematical Capability
The Reversal Curse [5], where LLMs trained from a sentence “A is B” are not able to generalize to answer “B is A”, also aligns with the observation in this paper that LLMs lack backward mathematical reasoning ability. To evaluate the backward mathematical capability, we propose a GSM8K-Backward test set, including 1270 backward questions by using SV and FOBAR to augment the original GSM8K test set (as shown in Example 3.3 and Example 3.3). Figure 7 shows the accuracy comparison of different 7B mathematical LLMs between the GSM8K and GSM8K-Backward datasets. As can be seen, existing LLMs struggle to solve mathematical problems in backward rationales and our MetaMath has a significant improvement on both datasets. Specifically, the ways where different LLMs solve the backward mathematical problem are illustrated through examples in Appendix C.
4.7 Reasoning Paths with Incorrect Answer Can Also Be Useful
Data Accuracy
GSM8K [13] 41.6
Incorrect Answers 43.6
Correct Answers 52.2
Table 4: Testing accuracy on GSM8K of LLaMA-2-7B trained on different data.
We conduct experiments on GSM8K using LLaMA-2-7B to study whether the answer augmentation samples with incorrect answers are helpful for finetuning the LLM. We randomly choose 7,473 reasoning paths with incorrect answers from the generated answers, and we ensure that the size is the same as that of the original training set. From Table 4, we observe that the model finetuned on the augmented data with incorrect answers is still better than SFT, which is counter-intuitive. We hypothesize that although the final answer is incorrect, some intermediate reasoning steps are correct (see Example 4.7). These reasoning steps can still be useful supervision signals. Our results are also aligned with [39], where they discover the importance of intermediate process supervision for reasoning.
Example 4.1:  A Reasoning Path with Incorrect Answer Question: Tonya is in a hamburger eating contest. Each hamburger is 4 ounces. Last year the winner ate 84 ounces. How many hamburgers does she have to eat to beat last year’s winner? (Ground-truth answer is $22$) Generated Answer: If each hamburger is 4 ounces and last year’s winner ate 84 ounces, then Tonya needs to eat 84/4 = 21 hamburgers to beat last year’s winner. The answer is: 21
4.8 More Data is not Always Better
@@
Example 4.2:  Case Study in Test Set of GSM8K Question: Darrell and Allen’s ages are in the ratio of 7:11, If their total age now is 162, calculate Allen’s age 10 years from now. (Ground-truth answer is $109$) SFT Answer: The total ratio representing their ages is 7+11=<<much-less-than$<$ <7+11=18>>much-greater-than$>$ >18. The fraction representing Darrell’s age is 7/18, and since their total age now is 162, Darrell’s age is 7/18*162 = <<much-less-than$<$ <7/18*162=60>>much-greater-than$>$ >60. The fraction representing Allen’s age is 11/18, and since their total age now is 162, Allen’s age is 11/18*162 = <<much-less-than$<$ <11/18*162=84>>much-greater-than$>$ >84. Ten years from now, Allen’s age will be 84 + 10 = <<much-less-than$<$ <84+10=94>>much-greater-than$>$ >94. RFT Answer: The total ratio representing Darrell and Allen’s ages is 7+11 = <<much-less-than$<$ <7+11=18>>much-greater-than$>$ >18. The fraction representing Darrell’s age from the total ratio is 7/18, and since their total age is 162, Darrell is 7/18*162 = 100 years old. Allen’s age, from the total ratio, is 11/18, and since Darrell is 100, Allen is 11/18*109 = <<much-less-than$<$ <11/18*100=66>>much-greater-than$>$ >66 years old. Ten years from now, Allen will be 66 + 10 = 76 years old. WizardMath Answer: We know Darrell and Allen’s ages are in the ratio of 7:11, which means the total ratio representing their ages is 7+11 = 18. If their total age now is 162, we can calculate Darrell’s age by using the ratio. Specifically, Darrell’s age is 7/18 of 162, which is 7/18*162 = 49 years old. … The answer is: 76. Our Answer: The ratio of Darrell’s age to Allen’s age is 7:11. Let’s assume Darrell’s age is 7x and Allen’s age is 11x. The total age of Darrell and Allen is 7x + 11x = 18x. We are given that the total age is 162, so 18x = 162. Dividing both sides by 18, we get x = 9. Therefore, Allen’s age is 11x = 11 * 9 = 99. 10 years from now, Allen’s age will be 99 + 10 = 109.
@@ 5 Concluding Remarks
In this paper, we focus on improving the mathematical problem-solving abilities of open-source LLMs. By bootstrapping mathematical questions on GSM8K and MATH, we present a high-quality and diverse dataset MetaMathQA, involving forward reasoning and backward reasoning samples. Our family of LLMs finetuned on MetaMathQA, called MetaMath, have achieved state-of-the-art on mathematical benchmarks among all open-source LLMs. Remarkably, MetaMath-7B reaches $66.5\%$ on GSM8K and $19.8\%$ on MATH, surpassing previous open-source LLMs by a significant margin. Our work further emphasizes the importance of the characteristics of the training data on boosting LLM problem-solving capabilities.
@@ Acknowledgement
This work was supported by NSFC key grant 62136005, NSFC general grant 62076118, and Shenzhen fundamental research program JCYJ20210324105000003. This research was supported in part by the Research Grants Council of the Hong Kong Special Administrative Region (Grants 16200021 and 16202523). AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, and the Leverhulme Trust via CFI.
Appendix
[…]
@@ A.1 Rephrasing Prompts
Example A.1:  Prompt for Rephrasing GSM8K Questions You are an AI assistant to help me rephrase questions. Follow the given examples. Question: Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left? Rephrase the above question: What is the amount of money that Olivia has left after purchasing five bagels for \$3 each, if she initially had \$23? Question: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday? Rephrase the above question: After losing 23 golf balls on Tuesday and an additional 2 on Wednesday, how many golf balls does Michael have left if he initially had 58 golf balls? Question: Angelo and Melanie want to plan how many hours over the next week they should study together for their test next week. They have 2 chapters of their textbook to study and 4 worksheets to memorize. They figure out that they should dedicate 3 hours to each chapter of their textbook and 1.5 hours for each worksheet. If they plan to study no more than 4 hours each day, how many days should they plan to study total over the next week if they take a 10-minute break every hour, include 3 10-minute snack breaks each day, and 30 minutes for lunch each day? Rephrase the above question: Angelo and Melanie need to study 2 chapters in their textbook and 4 worksheets for their upcoming test. They have planned to dedicate 3 hours for each chapter and 1.5 hours for each worksheet. They can study for a maximum of 4 hours each day, taking into account 10-minute breaks every hour, 3 10-minute snack breaks per day, and 30 minutes for lunch. How many days do they need to study in total over the next week to complete their study plan? Question: Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total? Rephrase the above question: If Leah had 32 chocolates and her sister had 42, and they both consumed 35 chocolates, what is the total number of chocolates that they have left? Question: There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room? Rephrase the above question: If there were initially nine computers in the server room and five more computers were added each day from Monday to Thursday, what is the current total number of computers in the server room? Question: Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny? Rephrase the above question: If Jason initially had 20 lollipops and now has 12 after giving some to Denny, how many lollipops did he give to Denny? Question: Sam bought a dozen boxes, each with 30 highlighter pens inside, for \$10 each box. He rearranged five of these boxes into packages of six highlighters each and sold them for \$3 per package. He sold the rest of the highlighters separately at the rate of three pens for \$2. How much profit did he make in total, in dollars? Rephrase the above question: Sam purchased 12 boxes, each containing 30 highlighter pens, at \$10 per box. He repackaged five of these boxes into sets of six highlighters and sold them for \$3 per set. He sold the remaining highlighters individually at a rate of three pens for \$2. What is the total profit he made in dollars? Question: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today? Rephrase the above question: If there were initially 15 trees in the grove and the grove workers are planning to plant more trees today, resulting in a total of 21 trees, how many trees did the workers plant today? Question: {Q} Rephrase the above question:
A.2 Rewriting Question with Answer into a Declarative Statement
Example A.2:  Prompts for Rewriting Question with Answer into a Declarative Statement You are an AI assistant to help me rewrite question into a declarative statement when its answer is provided. Follow the given examples and rewrite the question. Question: How many cars are in the parking lot? The answer is: 5. Result: There are 5 cars in the parking lot. Question: {Q} The answer is: {A}. Result:
Appendix B Experimental Details